# A Guide to Stochastic Optimisation for Large-Scale Inverse Problems

### v1

@@ 2       $x^{(k+1)}=\mathrm{prox}_{\tau g}\big{(}x^{(k)}-\tau\nabla h(x^{(k)})\big{)}$
Algorithm 1 Proximal Gradient Descent
@@ 4      $x^{(k+1)}=\mathrm{prox}_{\tau g}\big{(}y^{(k+1)}-\tau\nabla h(y^{(k+1)})\big{)}$
Algorithm 2 FISTA
@@ 4       $x^{(k+1)}=\mathrm{prox}_{\tau_{k}g}\big{(}x^{(k)}-\tau_{k}\tilde{\nabla}h(x^{(k)})\big{)}$
Algorithm 7 Stochastic Gradient Descent
@@ 3.2.3   Further Stochastic Gradient-Based Methods
Algorithms for the three-term sum from Section 2.2.2 also admit stochastic variants. For example, stochastic variants of PD3O or Condat–Vũ algorithms can be combined with SGD or SVRG in a mix-and-match fashion by replacing the gradient $\nabla h(x)$, with a corresponding stochastic estimator, see [SCMR22]. Using stochasticity in the dual variable is discussed in Section 3.3.
@@ 3.2.4   Accelerated Stochastic Variance-Reduced Gradient Methods
The first attempt along this direction is [Nit14], in which FISTA is combined with SVRG gradient estimator, however no theoretical guarantee when the problem is not strongly convex.
Later on in [LMH15], based on an accelerated proximal point algorithm, a so-called Catalyst acceleration is proposed to speeding up the solving of (24). Catalyst contains two loops, the outer loop is accelerated proximal point algorithm, while the inner loop is solved with a stochastic gradient scheme to certain accuracy. In [AZ17], an acceleration scheme named Katyusha is proposed for accelerating SAGA and SVRG. By incorporating a “negative momentum” term, the author manages to prove an $\mathcal{O}(1/k^{2})$ convergence rate without strong convexity. Later on in [DES22], a more elegant acceleration scheme is proposed, see Algorithm 10 below, and the authors show that the negative momentum is not needed to achieve the $\mathcal{O}(1/k^{2})$ convergence.
@@ 3       Sample $i_{k}\in\{1,\ldots,\ell\}$ uniformly at random
$A^{*}\overline{y}^{(k+1)}=(1+\ell)A^{*}_{i_{k}}(y^{(k+1)}_{i}-y^{(k)}_{i})+\sum_{i=1}^{\ell}A^{*}_{i}y^{(k)}_{i}.$
Note that this bears a close resemblance to a SAGA estimator of the gradient of the bilinear term in (27) with respect to $x$. However, due to the factor $1+\ell$ instead of $\ell$, this is a biased gradient estimator. Its effectiveness on large-scale problems has been shown in various applications such as PET with sinogram data [EMS19], PET with listmode data [SH22], CT [CDE+24, RDH21, PAD+21], multi-spectral CT [PAD+21], parallel MRI [GDE21, GDE22] and magnetic particle imaging [ZB21].
@@ 3.4   Further Reading
Standard second-order methods have seen limited application in large scale problems, due to the costs and intractability in computing the Hessian of the objective. As discussed in Section 2.3, deterministic quasi-Newton methods use information about the curvature without computing the Hessians. Instead, approximation techniques are used to approximate the Hessian or the corresponding matrix-vector products. A direct extension of quasi-Newton methods to the stochastic setting can be challenging, since noisy estimates of the curvature (coupled with stochastic estimators of the gradient) can destabilise the iterations. A number of methods have been proposed to solve this issue, and leverage the strengths of quasi-Newton and stochastic gradient methods.
A pioneering work in this field is the stochastic quasi-Newton method [SYG07], which introduced stochastic modifications of BFGS and L-BFGS, showing a sub-linear convergence rate. This limitation was addressed by [MNJ16], achieving linear convergence rates under mild assumptions, through a combination of the stochastic variant of L-BFGS [BHNS16] with variance reduction.
Sub-sampled quasi-Newton methods [RKDA14, EM15, XYR+16, RKM19] randomly sub-sample the Hessian to further reduce the costs, and use matrix concentration inequalities to ensure that information about the curvature is well preserved. Despite these advancements, stochastic quasi-Newton methods are still not widely adopted neither in the inverse imaging community, with notable counterexamples [MXC+19, HPAT22, PA21] and they continue to be an active area of research.
@@
are, in a certain sense, similar to the solutions of the original problem, but are drastically cheaper to compute.
Common sketching techniques include random projections, sub-sampling, and leverage score sampling [Woo14, DMM06, Mah11]. This type of sketching comes at a cost since there is an irreversible loss of information [PW16].
@@
An alternative way to perform sketching for inverse problems is via iterative Hessian sketching (IHS). As Newton’s method, it construct quadratic approximations to the original problem via Hessians which are then replaced by sketched versions of it. In more detail, consider minimising a smooth convex objective $h(x)$, at each iteration, a sketching matrix $S\in\mathbb{R}^{m\times d}$ is used to construct a sketched Hessian matrix $H_{S}=(H^{1/2})^{*}S^{*}SH^{1/2}$, where $H$ is the original Hessian matrix at the current iterate $x^{(k)}$ and $H^{1/2}$ is a square-root of it. The sketched Hessian can then be used as
@@
where $\tau_{k}>0$ is the step-size. For certain convex optimisation problem, IHS is faster than the standard scheme  [PW16, PW17]. It can also be combined with (accelerated) PGD [TGD17] and is applied to accelerate the multi-coil MRI reconstruction [OOI+24].
Particularly common are Gaussian sketches, which have the strongest theoretical guarantees (with respect to compression quality) but are computationally inefficient. They can be accelerated through the Johnson–Lindenstrauss transform (FJLT) [AC09, AL08], and the sparse Johnson–Lindenstrauss transform (SJLT) [CW13], to name a few techniques. In particular, SJLT trades off computational efficiency for theoretical guarantees, but it is often used in practice.
Although they have some clear advantages, sketching-based methods are rarely applied to inverse imaging problems. This can primarily be attributed to the fact that sketching may destroy any beneficial structure present in the forward operator. For example, in CT the Radon transform is a sparse matrix for which we can only use subsampling as the sketching, leading to suboptimal performance. The MRI operator includes FFTs which again prevent naive application of sketching. Recently in [OOI+24], a special sketching operator which can preserve the fast computational structure of FFT is developed and applied to multi-coil MRI reconstruction task.
@@
In the deep learning era, a significant advance of stochastic gradient descent in the context of training (deep) neural networks is the adaptive step-size stochastic gradient schemes, such as AdaGrad [DHS11], AdaDelta [Zei12], RMSProp [TH12], ADAM [KB15], and so on. We refer to [Rud16] for a brief overview of these methods.
The main motivation behind this type of methods is an observation, that during the training procedure, certain variables are updated much less compared to the others. More precisely, let $d^{(k)}=\tilde{\nabla}h(x^{(k)})$ be the stochastic gradient in each step, then up to $T$ step, if for some index $i$ such that $\sum\nolimits_{k=1}^{T}|d_{i}^{(k)}|$ is much smaller that the accumulation of the rest entries, this results in a much slower training behaviour of $x_{i}$, hence requires longer time for training. To overcome this problem, a basic idea is to set different step-sizes for each entries of $x$. To this end, let $D^{(T)}$ be the accumulation of the stochastic gradient $d^{(k)}$, for example for each $i=1,...,n$, let $D^{(T)}_{i}=\epsilon+\sum_{k=1}^{T}|d_{i}^{(k)}|$ where $\epsilon>0$ is for avoiding the sum is 00. Let $H_{T}=\mathrm{diag}(1/D^{(T)})$ be a diagonal matrix where the diagonal elements is the reverse of $D^{(T)}_{i}$, then an adaptive step-size stochastic gradient descent can take for following step
@@
Apparently, if for some $i$ the value $D^{(T)}_{i}$ is very small, then at current step $x^{(k+1)}_{i}$ will have a much larger step-size, hence receives larger update.
Compared to the standard SGD scheme, in general adaptive step-size schemes have two advantages: 1) make tuning of step-size automatic such that each entries converge at relatively same speed; 2) tuning-free step-size choices, for instance in (28) we only need to choose $\tau>0$. Over the past decade, adaptive step-size SGD schemes are more popular in training neural networks, while deeper theoretical understandings are obtained for the standard SGD.
@@
One thing we would like to highlight regarding adaptive step-size stochastic gradient schemes, is that this branch of algorithms is the most dominant choices in machine learning, especially for ADAM algorithm. In the field of inverse (imaging) problems, though the adoption of stochastic algorithms is quite trendy nowadays, adaptive step-size schemes seem received rather limited attention. In comparison, variance reduction schemes are more popular.
@@
To address this, new approaches have been developed in recent years that take the stochastic approach to the proximal domain. In [CR23] RandProx algorithm is introduced which uses a stochastic estimator for the proximal step, this refers to the prox of $f^{\ast}$) in the primal-dual Davis–Yin algorithm. The stochastic estimator is in its simplest form defined by a probability parameter $p$, i.e. depending on a coin flip the prox is either applied or not applied. ProxSkip [MMSR22] applies this methodology directly to proximal gradient-style algorithms. However, these methods are predominantly applied to federated or decentralised learning, and their use in inverse imaging problems has not yet been thoroughly investigated.
@@
For some suitably-structured nonlinear inverse problems, the stochastic gradient methods have been successfully applied to. For example the randomised Kaczmarz method and its SVRG-inspired variance reduced variants have been successfully applied to phase-retrieval problems [XLTW22] and EIT [vdDA12, ZQTL24]. However, in general the application of stochastic gradient methods can be more challenging than the inverse problems with a linear forward operators, due to the additional level of difficulty in the optimisation objective. One obvious potential issue of applying stochastic gradient methods on solving nonlinear inverse problem is the difficulty in performing back-tracking line search for step-sizes.
@@ 4   Challenges for Stochastic Optimisation for Inverse Problems
In the previous two sections, we have provided an introduction and overview of both deterministic and stochastic optimisation algorithms. In this section we present a thorough discussion on various theoretical and practical aspects of stochastic optimisation methods.
@@ 4.1   When and Why Are Stochastic Gradient Methods Fast?
As we mentioned in the beginning of Section 3, the aim of developing stochastic optimisation algorithms is to deal with large-scale optimisation problems. One may naturally wonder are stochastic optimisation methods guaranteed to be faster than their deterministic counterpart? Below are some key points.
@@
Whether an inverse problem is suitable or not for the application of stochastic gradient method largely depends on the structure of the optimisation problem. We start by introducing studies on fundamental linear inverse problems which can assist us to have understand better when and why stochastic gradient methods can be efficient/inefficient in different scenarios of inverse problems.
In the works of [TED19, TEGD20], a metric for estimating whether the inverse problem is suitable for applying stochastic gradient methods is proposed. This metric is named stochastic acceleration factor $\Upsilon:=L/L_{\max}$ which is proportional to the ratio of two Lipschitz constants. For the ideal cases, the ratio between the gradient Lipschitz constant is close to $n$. Many machine learning problems with common datasets fall into this category, as well as some important inverse imaging problems such as the CT imaging. However, there are many negative examples with $L_{\max}\approx L$, making $\Upsilon\approx 1$ where the stochastic gradient methods are not recommended. These include compressed sensing inverse problems, image denoising or image inpainting. For image deblurring it strongly depends on the (size of) the kernel, see also the numerical example in Section 5.
In the context of linear inverse problems where the data-fit is chosen to be the least-squares, a specialised analysis reveals that the success of stochastic acceleration is dependant on the structure of the forward operator. The theoretical analysis provided by [TEGD20] showed that an inverse problem is suitable for stochastic gradient methods if and only if the Hessian of the inverse problem has a fast-decaying spectrum. It is often neglected in the optimisation community, that stochastic gradient algorithms are in fact exploiting the intrinsic low-dimensional structure of the data matrices, and if such structure is weak in certain problems, then there will be no benefit using them.
@@
The importance of variance-reduced methods depends highly on the amount of noise in the measurement. It has been shown in [VML+19] that SGD converges linearly in the so-called interpolation regime, which in the context of inverse problems essentially means that there is no measurement noise. SGD can be even faster than variance-reduced methods as it can use a constant step-size of $1/(nL)$ in this case [GLQ+19]. In contrast variance-reduced methods typically require smaller constant step-sizes, for instance the largest known provably-convergent step-size is $1/(3nL)$ for SAGA [DBLJ14], while $1/(4nL)$ for SVRG [XZ14]. For the case where the measurement noise is nonnegligible, then the vanilla SGD must reduce its step-size for convergence and become slow for high accuracy solutions. In such cases, variance-reduced stochastic gradient methods are recommended due to their fast convergence in this regime which is prevalent for inverse problems. See also the numerical experiments in Section 5.
@@
As we discussed above, the stochastic acceleration factor test is a necessary condition for the theoretical usefulness of SGD for a specific inverse problem. Here we describe another test related to parallelisability. Practically, stochastic methods for inverse problems all split the forward operator $K$ into smaller blocks $K_{i}$ such that $(Kx)_{i}=K_{i}x$. A fundamentally important question is whether $K_{i}x$ for $i=1,\dots,s$ can be computed as fast as $Kx$. Stochastic methods essentially serialise the problem which adds a computational overhead. The practical feasibility of a stochastic method then depends on both the stochastic acceleration factor and the parallelisability of the forward operator. For example, most implementations of the X-ray transform are parallelisable over its angles and the forward operator in parallel MRI is parallelisable over the coils.
@@
The number of subsets allows us to interpolate between deterministic and high-variance random algorithms both of which have benefits and drawbacks. To illustrate this point we compare PGD with SGD. In every iteration of PGD, the most costly operations the evaluations of the gradient and the proximal operator. For imaging inverse problems often proximal operators with nonneglibable computational cost are used, e.g. the proximal operator of the total variation can be approximated via iterations [BT09, RC20] or a deep neural network applied [HLP21, TMTS24b]. In contrast, while SGD computes only a gradient on a subset, it does need to evaluate the proximal operator, too. Thus, an epoch of SGD has the same computational cost for the gradient, but has evaluated $n$-times the proximal operator which often leads to considerable overhead. Therefore, for most applications, the number of subsets $n$ needs to be chosen moderate to avoid this computational overhead.

### v2

@@ 2       $x^{(k+1)}=\mathrm{prox}_{\tau g}\big{(}x^{(k)}-\tau\nabla h(x^{(k)})\big{)}$
Algorithm 1 Proximal Gradient Descent (PGD)
@@ 4      $x^{(k+1)}=\mathrm{prox}_{\tau g}\big{(}y^{(k+1)}-\tau\nabla h(y^{(k+1)})\big{)}$
Algorithm 2 Fast Iterative Soft-shrinkage Algorithm (FISTA)
@@ 4       $x^{(k+1)}=\mathrm{prox}_{\tau_{k}g}\big{(}x^{(k)}-\tau_{k}\tilde{\nabla}h(x^{(k)})\big{)}$
Algorithm 7 Stochastic Gradient Descent (SGD)
@@ 3.2.3   Further Stochastic Gradient-Based Methods
Algorithms for the three-term sum from Section 2.2.2 also admit stochastic variants. For example, stochastic variants of PD3O or Condat–Vũ algorithms can be combined with SGD or SVRG in a mix-and-match fashion by replacing the gradient $\nabla h(x)$, with a corresponding stochastic estimator, see [SCMR22]. Similarly, for other first-order methods that contain gradient component, one can replace the gradient with a stochastic estimator, such as the primal-dual fixed-pint algorithm [ZZ21], the stochastic Davis–Yin three-operator splitting method [YVC16]. Notice that these stochastic methods use stochasticity to the primal variable, e.g. $\nabla h(x)$. While for dual variables, we leave the discussion in Section 3.3 as it leads to different type of stochastic optimisation methods.
@@ 3.2.4   Accelerated Stochastic Variance-Reduced Gradient Methods
The first attempt along this direction is [Nit14], in which FISTA is combined with SVRG gradient estimator, however no theoretical guarantee when the problem is not strongly convex. Later on in [LMH15], based on an accelerated proximal point algorithm, a so-called Catalyst acceleration is proposed to speeding up the solving of (24). Catalyst contains two loops, the outer loop is accelerated proximal point algorithm, while the inner loop is solved with a stochastic gradient scheme to certain accuracy. In [AZ17], an acceleration scheme named Katyusha is proposed for accelerating SAGA and SVRG. By incorporating a “negative momentum”, the author manages to prove an $\mathcal{O}(1/k^{2})$ convergence rate without strong convexity. In [DES22], a more elegant acceleration scheme is proposed, see Algorithm 10 below, and the authors show that the negative momentum is not needed to achieve the $\mathcal{O}(1/k^{2})$ convergence.
@@ 3       Sample $i_{k}\in\{1,\ldots,\ell\}$ uniformly at random
$A^{*}\overline{y}^{(k+1)}=(1+\ell)A^{*}_{i_{k}}(y^{(k+1)}_{i_{k}}-y^{(k)}_{i_{k}})+\sum_{i=1}^{\ell}A^{*}_{i}y^{(k)}_{i}.$
Note that this bears a close resemblance to a SAGA estimator of the gradient of the bilinear term in (27) with respect to $x$. However, due to the factor $1+\ell$ instead of $\ell$, this is a biased gradient estimator. Its effectiveness on large-scale problems has been shown in various applications such as PET with sinogram data [EMS19], PET with listmode data [SH22], PAT [RPDC20], CT [CDE+24, RDH21, PAD+21], multi-spectral CT [PAD+21], parallel MRI [GDE21, GDE22] and magnetic particle imaging [ZB21].
@@ 3.4   Further Reading
Standard second-order methods have seen limited application in large scale problems, due to the costs and intractability in computing the Hessian of the objective. As discussed in Section 2.3, deterministic quasi-Newton methods use information about the curvature without computing the Hessians. Instead, approximation techniques are used to approximate the Hessian or the corresponding matrix-vector products. A direct extension of quasi-Newton methods to the stochastic setting can be challenging, since noisy estimates of the curvature (coupled with stochastic estimators of the gradient) can destabilise the algorithm. A number of methods have been proposed to solve this issue, and leverage the strengths of quasi-Newton and stochastic gradient methods.
A pioneering work in this field is the stochastic quasi-Newton method [SYG07], which introduced stochastic modifications of BFGS and L-BFGS, showing a sub-linear convergence rate. This limitation was addressed by [MNJ16], achieving a linear convergence rate (under mild assumptions) through a combination of the stochastic variant of L-BFGS [BHNS16] with variance reduction.
Sub-sampled quasi-Newton methods, see e.g. [RKDA14, EM15, XYR+16, RKM19], randomly sub-sample the Hessian to further reduce the costs, and use matrix concentration inequalities to ensure that curvature information is well preserved. Despite these advances, stochastic quasi-Newton methods are still not widely adopted neither in the inverse imaging community, with notable counterexamples [MXC+19, HPAT22, PA21] and they continue to be an active research area.
@@
are, in a certain sense, similar to the solutions of the original problem but are drastically cheaper to compute.
Common sketching techniques include random projections, sub-sampling, and leverage score sampling [Woo14, DMM06, Mah11]. However, this type of sketching comes at a cost since there is an irreversible loss of information [PW16].
@@
An alternative approach for solving inverse problems via sketching is through iterative Hessian sketching (IHS). Similarly to Newton’s method, IHS constructs quadratic approximations to the original problem using Hessians, which are then sketched. In more detail, consider the minimisation of a smooth convex objective $h(x)$. At each iteration, a sketching matrix $S\in\mathbb{R}^{s\times d}$ is used to construct a sketched Hessian matrix $H_{S}=(H^{1/2})^{*}S^{*}SH^{1/2}$, where $H$ is the Hessian matrix at the current iterate $x^{(k)}$. IHS then follows the iterations
@@
where $\tau_{k}>0$ is the step-size. For certain convex optimisation problems IHS is faster than baseline methods  [PW16, PW17]. It can also be combined with (accelerated) PGD [TGD17], and it was used to accelerate multi-coil MRI reconstruction [OOI+24].
A particularly common choice for the sketching matrix are Gaussian sketches, which have the strongest theoretical guarantees (with respect to compression quality) but are computationally inefficient. They can be accelerated through the Johnson–Lindenstrauss transform (FJLT) [AC09, AL08], and the sparse Johnson–Lindenstrauss transform (SJLT) [CW13], to name a few techniques. In particular, SJLT trades off computational efficiency for theoretical guarantees, but it is often used in practice.
Although they have some clear advantages, sketching-based methods are rarely applied to inverse imaging problems. This can primarily be attributed to the fact that sketching may destroy any beneficial structure present in the forward operator. For example, in CT the Radon transform is a sparse matrix for which we can only use subsampling as the sketching, leading to suboptimal performance. The MRI operator includes FFTs which again prevents naive application of sketching. Recently in [OOI+24], a special sketching operator which can preserve the fast computational structure of FFT is developed and applied to multi-coil MRI reconstruction task.
@@
A significant advancement of SGD in the context of training (deep) neural networks are adaptive step-size stochastic gradient schemes, such as AdaGrad [DHS11], AdaDelta [Zei12], RMSProp [TH12], ADAM [KB15], to name a few. We refer to [Rud16] for a brief overview of these methods.
The main motivation behind this type of methods is the observation that some variables, or components of the solution, are updated significantly less during the training than other variables. More precisely, let $d^{(k)}=\tilde{\nabla}h(x^{(k)})$ denote the stochastic gradient computed in $k$-th iteration, for $x^{(k)}\in\mathbb{R}^{d}$. Then assume that for an iteration $T$ there exists an $1\leq i\leq d$, such that the accumulation of update directions, defined as $D_{i}^{(T)}=\sum\nolimits_{k=1}^{T}|d_{i}^{(k)}|$ is small compared to $D_{j}^{(T)}$ for $i\neq j$. In this case the corresponding entry $x_{i}$ of the solution will require a significantly longer training time. Adaptive steps-size methods overcome this issue by using step-sizes tailored to each entry of $x$. To this end, let $D^{(T)}$ be the accumulation of the stochastic gradient $d^{(k)}$. For example, as above let $D^{(T)}_{i}=\epsilon+\sum_{k=1}^{T}|d_{i}^{(k)}|$, for $i=1\ldots,n$, and a small positive constant $\epsilon>0$. For a diagonal matrix $H_{T}=\mathrm{diag}(1/D_{i}^{(T)})$ the adaptive step-size SGD can follow the iterations
@@
Therefore, if for some $i$ the value $D^{(k)}_{i}$ is very small, then the corresponding entry of $H_{k}$ will be large, and hence the entry $x^{(k+1)}_{i}$ will use a larger effective step-size and receive a larger update.
Compared to the standard SGD scheme, in general adaptive step-size schemes have two advantages. First, tuning of step-size becomes automatic such that each entries converge at relatively same speed. Second, step-size choices are tuning-free, for instance in (28) we only need to choose $\tau>0$. Over the past decade, adaptive step-size SGD schemes have been more popular in training neural networks, while deeper theoretical understandings are obtained for the standard SGD.
@@
We would like to highlight that adaptive step-size stochastic gradient schemes are the dominant optimisation schemes in machine learning, especially the ADAM algorithm. In the field of inverse (imaging) problems, even though the adoption of stochastic algorithms is quite trendy nowadays, adaptive step-size schemes have received rather limited attention. In comparison, variance reduction schemes are more popular.
@@
To address this, new approaches have been developed in recent years that take the stochastic approach to the proximal domain. In [CR23] RandProx algorithm is introduced which uses a stochastic estimator for the proximal step in the primal-dual Davis–Yin algorithm. The stochastic estimator is in its simplest form defined by a probability parameter $p$, i.e. depending on a coin flip the proximal operator is either applied or not applied. ProxSkip [MMSR22] applies this methodology directly to proximal gradient-style algorithms. However, these methods are predominantly applied to federated or decentralised learning, and their use in inverse imaging problems has not yet been thoroughly investigated.
@@
Stochastic gradient methods have been successfully applied to some suitably-structured nonlinear inverse problems. For example the randomised Kaczmarz method and its SVRG-inspired variance reduced variants have been successfully applied to phase-retrieval problems [XLTW22], EIT [vdDA12, ZQTL24], optical tomography [MAP20], and other problems. However, in general the application of stochastic gradient methods for non-linear problems poses additional challenges, due to the complexities introduced by the forward model and its properties. Another obvious potential issue in applying stochastic gradient methods for solving nonlinear inverse problem is the difficulty in performing back-tracking line search for step-sizes. These issues often require a problem specific solution and more research is needed in this area.
@@ 4   Challenges for Stochastic Optimisation for Inverse Problems
In the previous two sections we provided an overview of both deterministic and stochastic optimisation algorithms. In this section we present a thorough discussion on various theoretical and practical aspects of stochastic optimisation methods.
@@ 4.1   When and Why Are Stochastic Gradient Methods Fast?
As we mentioned in the beginning of Section 3, the aim of developing stochastic optimisation algorithms is to deal with large-scale optimisation problems. One may naturally wonder whether stochastic optimisation methods are guaranteed to be faster than their deterministic counterparts. Below are some key points.
@@
Whether an inverse problem is suitable or not for the application of a stochastic gradient method largely depends on the structure of the optimisation problem. which can assist us in better understanding when and why stochastic gradient methods can be efficient or inefficient in different inverse problems scenarios.
In the works of [TED19, TEGD20], a metric for estimating whether the inverse problem is suitable for applying stochastic gradient methods is proposed. This metric, named the stochastic acceleration factor, is defined as the ratio of the two Lipschitz constants $\Upsilon:=L/L_{\max}$. In the ideal cases for stochastic methods, the ratio between the gradient Lipschitz constant is close to $n$. Many machine learning problems with common datasets fall into this category, as well as some important inverse imaging problems such as CT reconstruction. However, there are many negative examples with $L_{\max}\approx L$, making $\Upsilon\approx 1$, in which case stochastic gradient methods are not recommended. These include compressed sensing inverse problems, image denoising or image inpainting. For image deblurring it strongly depends on the size of the kernel, see also the numerical example in Section 5.
In the context of linear inverse problems where the data-fit is given by the least-squares, a specialised analysis reveals that the success of stochastic acceleration depends on the structure of the forward operator. Theoretical analysis provided by [TEGD20] shows that an inverse problem is suitable for stochastic gradient methods if and only if the Hessian of the inverse problem has a fast-decaying spectrum. In the optimisation community, it is often overlooked that stochastic gradient methods are exploiting the intrinsic low-dimensional structure of the data matrices. If such a structure is weak in certain problems, the use of stochastic algorithms may not result in any benefits.
@@
The importance of variance-reduced methods depends highly on the amount of noise in the measurements. It has been shown in [VML+19] that SGD converges linearly in the so-called interpolation regime, which in the context of inverse problems essentially equates to a noise-free measurement regime. In this case SGD can even be faster than variance-reduced methods as it can use a constant step-size of $1/(nL_{\max})$[GLQ+19]. In contrast, variance-reduced methods typically require smaller constant step-sizes. For instance, the largest known provably-convergent step-size is $1/(3nL_{\max})$ for SAGA [DBLJ14], and $1/(4nL_{\max})$ for SVRG [XZ14]. When the measurement noise is not negligible vanilla SGD requires using a decaying step-size, in order to ensure convergence, and becomes slow for high accuracy solutions. In such cases, variance-reduced stochastic gradient methods are recommended due to their fast convergence in this regime which is prevalent for inverse problems. See also the numerical experiments in Section 5.
@@
As we discussed above, testing the stochastic acceleration factor is a necessary condition for the theoretical utility of SGD for specific inverse problems. Here we describe another test related to parallelisability. Practically, stochastic methods for inverse problems all partition the forward operator $K$ into smaller blocks $K_{i}$ such that $(Kx)_{i}=K_{i}x$. A fundamentally important question is whether $K_{i}x$,for $i=1,\dots,s$ can be computed as fast, or faster than $Kx$. Stochastic methods essentially serialise the problem which adds a computational overhead. The practical feasibility of a stochastic method then depends on both the stochastic acceleration factor and the parallelisability of the forward operator. For example, most implementations of the X-ray transform are parallelisable over its angles and the forward operator in parallel MRI is parallelisable over the coils.
@@
The number of subsets allows us interpolating between deterministic and high-variance stochastic algorithms, both of which have their benefits and drawbacks. To illustrate this point we compare PGD with SGD. In every iteration of PGD, the most costly operations are the evaluation of the gradient and of the proximal operator. In imaging inverse problems it is common to use proximal operators with a non-negligible computational cost, e.g. the proximal operator of the total variation, which can be approximated via iterations [BT09, RC20] or by applying a deep neural network [HLP21, TMTS24b]. Thus, a single data pass of SGD has the same computational cost for the gradient (as it is evaluated only on a subset), but it requires evaluating the proximal operator $n$ times, which often leads to a considerable overhead. Therefore, for most applications, the number of subsets $n$ needs to be chosen conservatively to avoid this computational overhead.